OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at (11, 7) and a radius of 13? (x – 11)2 + (y – 7)2 = 169 (x – 11)2 + (y + 7)2 = 169 (x + 11)2 + (y – 7)2 = 169 (x + 11)2 + (y + 7)2 = 169

OpenStudy (sirm3d):

i'll give the equation in "general" form, given its center (h,k) and radius r. \[\Large (x-h)^2+(y-k)^2 = r^2\]

OpenStudy (phi):

Check this http://www.regentsprep.org/Regents/math/algtrig/ATC1/circlelesson.htm but in general, if you replace the center (11,7) you should get 0 , 0

OpenStudy (phi):

I mean replace x with 11 in the equation and y with 7 in the equation if you get 0 for both, then 11,7 is the center

OpenStudy (sirm3d):

which one do you think is the correct answer, @brittneyy18rm ?

OpenStudy (anonymous):

still a little confused :/

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\Large (x-h)^2+(y-k)^2 = r^2\] is the general formula for a circle with center \((h,k)\) and radius \(r\) in your case you have the center is \((11, 7)\) and the radius is 13 so \(h=11,k=7,r=13\)

OpenStudy (anonymous):

(x – 11)2 + (y + 7)2 = 169 ???

OpenStudy (anonymous):

make a direct substitution onto the formula

OpenStudy (anonymous):

would the answer be the second choice then?

OpenStudy (anonymous):

should be a minus sign between the y and the 7

OpenStudy (anonymous):

the minus sign lives in the formula, so you should have \[(x-11)^2+(y-7)^2=13^2\]

OpenStudy (anonymous):

my bad, so it would be my first choice?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I get it now :)